OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y - x2 = -11x + 24? Answer (-3, 0) and (8, 0) (3, 0) and (-8, 0) (-3, 0) and (-8, 0) (3, 0) and (8, 0)

OpenStudy (anonymous):

you need to solve the equation x^2 - 11x + 24 = 0 can u do this?

OpenStudy (anonymous):

this can be factored find 2 numbers with the product = +24 and their sum = -11

OpenStudy (anonymous):

y - x²= -11x + 24 x intercept means set y = 0 y = x² -11x + 24

OpenStudy (anonymous):

Maria, give it a try!